# A Multi-Stage Framework with Taxonomy-Guided Reasoning for Occupation Classification Using Large Language Models

### v1

@@ Abstract
Automatically annotating job data with standardized occupations from taxonomies, known as occupation classification, is crucial for labor market analysis. However, this task is often hindered by data scarcity and the challenges of manual annotations. While large language models (LLMs) hold promise due to their extensive world knowledge and in-context learning capabilities, their effectiveness depends on their knowledge of occupational taxonomies, which remains unclear. In this study, we assess the ability of LLMs to generate precise taxonomic entities from taxonomy, highlighting their limitations. To address these challenges, we propose a multi-stage framework consisting of inference, retrieval, and reranking stages, which integrates taxonomy-guided reasoning examples to enhance performance by aligning outputs with taxonomic knowledge. Evaluations on a large-scale dataset show significant improvements in classification accuracy. Furthermore, we demonstrate the framework’s adaptability for multi-label skill classification. Our results indicate that the framework outperforms existing LLM-based methods, offering a practical and scalable solution for occupation classification and related tasks across LLMs.
@@ 1 Introduction
However, automatic occupation classification is inherently challenging due to the scarcity of large publicly available labeled datasets which require laborious efforts of labor market experts. The lack of data resources further hinders the development and evaluation of robust supervised learning-based classification models. Given the sheer size and complexity of occupational taxonomies. Non-expert annotators would require rigorous training to effectively interpret and apply the nuanced structures of these classifications, further complicating the issues.
Recent advancements in large language models (LLMs) have shown promises for addressing these challenges through prompt engineering (Li, Kang, and De Bie 2023; D’Oosterlinck et al. 2024). Despite their potential, critical research gaps remain unaddressed. First, the extent to which LLMs rely on their knowledge of occupational taxonomies for classification has not been explored. LLMs are pre-trained on broad datasets, which may include partial representations of occupational taxonomies. The quality and coverage of this taxonomic knowledge can also vary significantly across LLMs due to differences in their training data and architectures. Consequently, their classification effectiveness may be inconsistent. This raises important questions about the reliability and accuracy of LLM-based methods, particularly when dealing with less common or highly specific occupational categories, where the knowledge gaps are likely to be most pronounced.
@@
To address these research gaps, we conduct a comprehensive knowledge assessment study to evaluate the capabilities of various LLMs in generating accurate SOC titles and codes from the ONET-SOC taxonomy. Building on these findings, we propose a lightweight multi-stage framework that integrates inference, retrieval, and reranking components. Unlike computationally intensive fine-tuning or complex model modifications, our framework relies on efficient prompting and retrieval strategies, making it adaptable to a variety of LLMs without requiring additional training. By leveraging the detailed class information embedded in ONET-SOC, this framework enhances the accuracy and robustness of occupation classification tasks. Finally, we evaluate the framework on both occupation and skill classification tasks, demonstrating its effectiveness and adaptability across different labor-market domains.
Our contributions are threefold: (1) We are the first to systematically investigate LLM knowledge of the O*NET-SOC taxonomy through a series of generation tasks, providing insights into their strengths and limitations. (2) We propose a simple yet effective method to integrate grounded rationales into in-context examples, aligning LLM reasoning and outputs with taxonomic information while requiring minimal prompt engineering. (3) We explicitly examine the relationship between LLMs’ taxonomic knowledge and their downstream classification performance. Our findings demonstrate that leveraging external taxonomic information throughout the classification pipeline, from inference to reranking, enhances alignment with the taxonomy and improves overall performance across different LLMs.
@@ 3.3 Task Design
We devise a set of generation tasks to assess LLM knowledge by instructing LLMs to generate SOC titles and codes in order to assess both the ability to directly retrieve taxonomy information and the capacity to understand its hierarchical structure at varying levels of granularity.
@@
Additionally, we focus on two main context-to-output generation directions: code-to-title (CT) and title-to-code (TC). In CT, we test the model’s ability to generate a title (output) from a given code (context), while in the TC direction, we test its ability to generate a code (output) from a given title (context). To assess hierarchical understanding, each task direction is examined at two levels of granularity: 8-digit and 2-digit codes. These combinations yield three task variants: CT8 (title generation from 8-digit codes), TC8 (8-digit code generation from titles), and TC2 (2-digit code generation from titles). We exclude CT2 (title generation from 2-digit codes) as multiple titles can correspond to one 2-digit code.
@@ 3.4 Research Questions
RQ2: To what extent do LLMs internalize the hierarchical relationships within the O*NET-SOC taxonomy, demonstrated by their performance across different levels of granularity?
@@ 3.6 Metrics and Statistical Analysis
We evaluate performance using accuracy, defined as the proportion of generated answers that exactly match the corresponding values in the O*NET-SOC taxonomy. Statistical analyses, including a one-way ANOVA and linear regression, were performed using Python’s SciPy (1.11.3) and Statsmodels (0.14) libraries. The data used in the analyses meet the normality assumption.
@@ 3.7 Knowledge Assessment Results
RQ2: To what extent do LLMs internalize the hierarchical relationships within the O*NET-SOC taxonomy, demonstrated by their performance across different levels of granularity?
To investigate the impact of hierarchical granularity on model performance, a one-way ANOVA was performed to compare LLM accuracy on tasks involving 8-digit (specific occupations) and 2-digit (major occupational groups) SOC codes. The results demonstrate a statistically significant difference between the two levels ($F=12.667$, $p<0.001$), with LLMs performing significantly better on 2-digit tasks. The findings suggest that while LLMs exhibit some understanding of the hierarchical structure of the O*NET-SOC taxonomy, their ability to generate precise occupational titles and codes declines as the hierarchical granularity increases.
@@
Altogether, these findings suggest that while current LLMs exhibit varying degrees of knowledge regarding the O*NET-SOC taxonomy, their ability to fully internalize and apply hierarchical understanding of the taxonomy is limited, except for GPT-4o. Task design elements, particularly code granularity and output type, should be considered when leveraging LLMs for downstream applications like occupation classification.
@@ 5.2 Experimental Setup
LLMs: Each method was tested on five efficient LLMs: GPT-3.5 Turbo, Gemini-1.5 Flash, Claude Instant 1.2 , Llama 3 8B Instruct, and Mistral Small. As previously mentioned, these LLMs were chosen for their cost effectiveness, which make them practical for large-scale classification tasks. GPT-3.5 Turbo was accessed in June 2024, while the remaining models were accessed in October 2024. The temperature was set to zero for all models to guarantee consistent results.
@@
All prompt templates and in-context examples are included in Section B.2 of the appendix.
@@ 5.3 Evaluation Metric
To evaluate the performance of different methods, we used Precision@K ($K=1$), defined as the proportion of correctly classified labels among the top-K predicted labels (SOC titles and 8-digit codes). The choice of $K=1$ is appropriate as only the top-predicted label is typically actionable in occupation classification.
Since ground-truth labels for the job records in Jobs12K were not available, we employed GPT-4o, which has been shown to be highly knowledgeable of the O*NET-SOC taxonomy in Section 3, to identify the most relevant label. Specifically, we prompted GPT-4o to select the most appropriate SOC label from the pool of predicted labels generated by all tested methods for each job record. To mitigate positional bias, the lists of candidate labels were shuffled. GPT-4o was also instructed to explicitly indicate if none of the candidates were applicable. This evaluation was repeated for 3 runs, and majority voting was applied to select the final labels. In the case of ties, the final selection was made randomly. The prompt template for occupation selection is included in Figure 7 in Section B.2.
@@ 5.4 Results
In summary, the findings demonstrate the effectiveness of the proposed framework, which combines LLM reasoning capabilities with taxonomy-guided examples, retrieval, and reranking to achieve robust performance across efficient LLMs in occupation classification tasks.
@@ 6.2 Experimental Setup
We adopted the same experimental setup as described in Section 5.2, including LLM models and temperature setting, reasoning methods, retrieval units, and ablation variants. The LLMs were also accessed during the same periods as described in the occupation classification evaluation. All prompt templates and in-context examples are provided in Section B.3 of the appendix.
To handle the extensive label space in the ESCO taxonomy, we set the number of top candidates retrieved during the retrieval stage to $m=50$, following D’Oosterlinck et al. (D’Oosterlinck et al. 2024). In the reranking stage, the number of top-$n$ candidates was set to $n=10$. The number of in-context examples for both TGRE and CoT reasoning methods during inference and reranking was empirically set to $k=2$, as determined by the optimization module used in IReRa (D’Oosterlinck et al. 2024).
@@ 6.4 Results
Table 5 presents the results for RP@10 across methods and models. Due to space limitations, RP@3 and RP@5 results are provided in Tables 8 and  9 in Section  C of the appendix. The best performance is achieved by TGRE* with sentence-based retrieval, attaining an RP@10 of 0.7052 on Gemini-1.5 Flash, outperforming IReRa (0.6546) [^2] Note that the state-of-the-art RP@10 of 0.7058 was achieved by the best IReRa variant as reported in their original experiment (D’Oosterlinck et al. 2024) though we were unable to independently replicate this result. and LLM4Jobs (0.6151). CoT* also greatly outperforms the best CoT variant across all models. Consistent with previous findings (Clavié and Soulié 2023; D’Oosterlinck et al. 2024), these results reinforce the effectiveness of GPT-4 as a powerful reranker for skill classification.
@@ 8 Conclusion and Implications
In this paper, we investigate the capabilities of large language models (LLMs) in understanding the ONET-SOC taxonomy. The study begins with a knowledge assessment experiment to evaluate the performance of both frontier and efficient LLMs in generating precise occupational titles and codes. The results indicate that frontier models, especially GPT-4o, consistently outperform the efficient ones, demonstrating their in-depth taxonomic understanding. However, LLMs struggle to fully understand the taxonomy’s hierarchical structure, as evidenced by their lower performance in generating specific occupations compared to major occupations.
@@ Appendix A The Lightcast Dataset
Appendix B Prompt Templates and In-context Examples
@@ B.1 Knowledge Assessment
B.2 Occupation Classification
This section describes the prompt templates and in-context examples used in the inference and reranking stages of our framework for the occupation classification task. Figures  3 and  4 display TGRE-based and CoT-based prompt templates for occupation classification, respectively. Figures  5 and  6 display k-shot in-context examples for TGRE-based and CoT-based prompt templates for occupation classification, respectively.

### v2

@@ Abstract
Automatically annotating job data with standardized occupations from taxonomies, known as occupation classification, is crucial for labor market analysis. However, this task is often hindered by data scarcity and the challenges of manual annotations. While large language models (LLMs) hold promise due to their extensive world knowledge and in-context learning capabilities, their effectiveness depends on their knowledge of occupational taxonomies, which remains unclear. In this study, we assess the ability of LLMs to generate precise taxonomic entities from taxonomy, highlighting their limitations, especially for smaller models. To address these challenges, we propose a multi-stage framework consisting of inference, retrieval, and reranking stages, which integrates taxonomy-guided reasoning examples to enhance performance by aligning outputs with taxonomic knowledge. Evaluations on a large-scale dataset show that our framework not only enhances occupation and skill classification tasks, but also provides a cost-effective alternative to frontier models like GPT-4o, significantly reducing computational costs while maintaining strong performance. This makes it a practical and scalable solution for occupation classification and related tasks across LLMs.
@@ 1 Introduction
However, automatic occupation classification is inherently challenging due to the scarcity of large publicly available labeled datasets which require laborious efforts of labor market experts. The lack of data resources further hinders the development and evaluation of robust supervised learning-based classification models. Given the sheer size and complexity of occupational taxonomies, non-expert annotators would require rigorous training to effectively interpret and apply the nuanced structures of these classifications, further complicating the issues.
Recent advancements in large language models (LLMs) have shown promises for addressing these challenges through prompt engineering (Li, Kang, and De Bie 2023; D’Oosterlinck et al. 2024). Despite their potential, critical research gaps remain unaddressed. First, the extent to which LLMs rely on their knowledge of occupational taxonomies for classification has not been explored. LLMs are pre-trained on broad datasets, which may include partial representations of occupational taxonomies. The quality and coverage of this taxonomic knowledge can also vary significantly across LLMs due to differences in their training data, architectures, and model sizes. Consequently, their classification effectiveness may be inconsistent. This raises important questions about the reliability and accuracy of LLM-based methods, particularly when dealing with less common or highly specific occupational categories, where the knowledge gaps are likely to be most pronounced.
@@
To address these research gaps, we conduct a knowledge assessment study to evaluate the capabilities of various LLMs in generating accurate SOC titles and codes from the ONET-SOC taxonomy. Our findings reveal two major shortcomings: (1) LLMs, particularly smaller models, often struggle with generating precise SOC codes, particularly for less common occupations, and (2) their classification accuracy varies significantly based on hierarchical granularity, indicating limited internalization of taxonomic knowledge. Building on these findings, we propose a multi-stage framework that integrates inference, retrieval, and reranking components. Specifically, we incorporate external knowledge from the O*NET taxonomy to guide LLM’s reasoning through grounded in-context examples. Unlike computationally intensive fine-tuning or complex model modifications, our framework relies on efficient prompting and retrieval strategies, making it adaptable to a variety of LLMs without requiring additional training. By leveraging the detailed class information embedded in ONET-SOC, this framework enhances the accuracy and robustness of occupation classification tasks. Finally, we evaluate the framework on both occupation and skill classification tasks, demonstrating its effectiveness in mitigating the knowledge gaps and its adaptability across domains.
Our contributions are fourfold:
  1. 1.
    We are the first to systematically investigate LLM knowledge of the O*NET-SOC taxonomy through a series of generation tasks, providing insights into their strengths and limitations.
  2. 2.
    We propose a novel framework that builds on retrieval-based prompting methods (Li, Kang, and De Bie 2023; D’Oosterlinck et al. 2024) by systematically integrating grounded rationals into in-context examples, aligning LLM reasoning and outputs with taxonomic information while minimizing prompt engineering effort.
  3. 3.
    We explicitly examine the relationship between LLMs’ taxonomic knowledge and their downstream classification performance. Our findings demonstrate that leveraging external taxonomic information throughout the classification pipeline, from inference to reranking, enhances alignment with the taxonomy and improves overall performance across different LLMs.
  4. 4.
    We demonstrate that our framework significantly reduces the computational costs while maintaining high classification accuracy. By enabling smaller, more efficient LLMs to achieve performance comparable to frontier models like GPT-4o, the framework provides a scalable and cost-effective alternative to GPT-4o’s direct prompting, making it practical for large-scale classification tasks.
@@ 3.3 Task Design
We devise a set of generation tasks to assess LLM knowledge by instructing LLMs to generate SOC titles and codes in order to assess both the ability to directly retrieve taxonomy information and the capacity to perform at varying levels of granularity within the taxonomy.
@@
Additionally, we focus on two main context-to-output generation directions: code-to-title (CT) and title-to-code (TC). In CT, we test the model’s ability to generate a title (output) from a given code (context), while in the TC direction, we test its ability to generate a code (output) from a given title (context). To analyze model performance across varying levels of granularity, each task direction is examined using both 8-digit and 2-digit codes. These combinations yield three task variants: CT8 (title generation from 8-digit codes), TC8 (8-digit code generation from titles), and TC2 (2-digit code generation from titles). We exclude CT2 (title generation from 2-digit codes) as multiple titles can correspond to one 2-digit code.
@@ 3.4 Research Questions
RQ2: How does LLM performance vary across different levels of granularity within the O*NET-SOC taxonomy?
@@ 3.6 Metrics and Statistical Analysis
We evaluate performance using accuracy, defined as the proportion of generated answers that exactly match the corresponding values in O*NET-SOC. The average accuracy score provides an overall metric measuring the model’s ability to generate taxonomic entities across tasks. A one-way ANOVA and linear regression were performed using Python’s SciPy (1.11.3) and Statsmodels (0.14) libraries. The data used in the analyses satisfy normality assumption.
@@ 3.7 Knowledge Assessment Results
RQ2: How does LLM performance vary across different levels of granularity within the O*NET-SOC taxonomy?
To analyze how LLM performance varies across different levels of granularity, a one-way ANOVA was performed to compare LLM accuracy on tasks involving 8-digit (specific occupations) and 2-digit (major occupational groups) SOC codes. The results demonstrate a statistically significant difference between the two levels ($F=12.667$, $p<0.001$), with LLMs performing significantly better on 2-digit tasks. The findings suggest that LLM performance varies with the granularity of the occupational codes. This disparity may be influenced by several factors, including the inherent difficulty of predicting finer-grained categories and potential limitations in the LLMs’ ability to fully leverage hierarchical information within the taxonomy.
@@
Altogether, these findings suggest that while current LLMs exhibit varying degrees of knowledge regarding the O*NET-SOC taxonomy, their performance in tasks involving the taxonomy varies, with GPT-4o demonstrating the strongest performance. Factors such as task design elements, particularly code granularity and output type, should be considered when leveraging LLMs for downstream applications like occupation classification.
@@ 5.2 Experimental Setup
LLMs: Each method was tested on five efficient LLMs: GPT-3.5 Turbo, Gemini-1.5 Flash, Claude Instant 1.2 , Llama 3 8B Instruct, and Mistral Small. As previously mentioned, these LLMs were chosen for their cost effectiveness, which make them practical for large-scale classification tasks. GPT-3.5 Turbo was accessed in June 2024, while the remaining models were accessed in October 2024. The temperature was set to zero for all tasks and LLMs to reduce randomness and promote deterministic outputs.
@@
All prompt templates and in-context examples are included in Section D.2 of the appendix.
@@ 5.3 Evaluation Metric
We used Precision@K ($K=1$), defined as the proportion of correctly classified labels among the top-K predicted labels (SOC titles and 8-digit codes). $K=1$ is appropriate as occupation classification typically considers only the top prediction actionable. As $K$=1 and each instance has a single ground-truth label, Precision@1 is equivalent to Recall@1, F1@1, and Accuracy@1.
Since ground-truth labels for the job records in Jobs12K were not available, we employed GPT-4o, which has been shown to be highly knowledgeable of the O*NET-SOC taxonomy in Section 3, to identify the most relevant label. Specifically, we prompted GPT-4o to select the most appropriate SOC label from the pool of predicted labels generated by all tested methods for each job record. To mitigate positional bias, the lists of candidate labels were shuffled. GPT-4o was also instructed to explicitly indicate if none of the candidates were applicable. This evaluation was repeated for 3 runs, and majority voting was applied to select the final labels. In the case of ties, the final selection was made randomly. The prompt template for occupation selection is included in Figure 7 in Section D.2.
@@ 5.4 Results
In summary, the findings demonstrate the effectiveness of the proposed framework, which combines LLM reasoning capabilities with taxonomy-guided examples, retrieval, and reranking to achieve robust performance across efficient LLMs in occupation classification tasks. Beyond performance improvements, the computational cost analysis in Appendix G indicates that the framework is highly cost-effective. While the framework incurs marginal processing time due to retrieval and reranking, our analysis shows that it substantially reduces API costs by 72% compared to direct GPT-4o inference, making it a more practical and scalable approach for large-scale classification tasks.
@@ 6.2 Experimental Setup
We adopted the same experimental setup as described in Section 5.2, including LLM models and temperature setting, reasoning methods, retrieval units, and ablation variants. The LLMs were also accessed during the same periods as described in the occupation classification evaluation. All prompt templates and in-context examples are provided in Section D.3 of the appendix.
To handle the extensive label space in the ESCO taxonomy, we set the number of top candidates retrieved during the retrieval stage to $m=50$, following D’Oosterlinck et al. (D’Oosterlinck et al. 2024). In the reranking stage, the number of top-$n$ candidates was set to $n=10$. The number of in-context examples for both TGRE and CoT reasoning methods during inference and reranking was set to $k=2$. For CoT, these in-context examples were selected by the optimization module in IReRa (D’Oosterlinck et al. 2024), whereas they were randomly chosen in TGRE-based variants.
@@ 6.4 Results
Table 5 presents the results for RP@10 across methods and models. Due to space limitations, RP@3 and RP@5 results are provided in Tables 8 and  9 in Section  E of the appendix. The best performance is achieved by TGRE* with sentence-based retrieval, attaining an RP@10 of 0.7052 on Gemini-1.5 Flash, outperforming IReRa (0.6546) [^2] Note that the state-of-the-art RP@10 of 0.7058 was achieved by the best IReRa variant as reported in their original experiment (D’Oosterlinck et al. 2024) though we were unable to independently replicate this result. and LLM4Jobs (0.6151). CoT* also greatly outperforms the best CoT variant across all models. Consistent with previous findings (Clavié and Soulié 2023; D’Oosterlinck et al. 2024), these results reinforce the effectiveness of GPT-4 as a powerful reranker for skill classification.
@@ 8 Conclusion and Implications
In this paper, we investigate the capabilities of large language models (LLMs) in understanding the ONET-SOC taxonomy. The study begins with a knowledge assessment experiment to evaluate the performance of both frontier and efficient LLMs in generating precise occupational titles and codes. The results indicate that frontier models, especially GPT-4o, consistently outperform the efficient ones, demonstrating their in-depth taxonomic understanding. However, LLM performance varies across different levels of code granularity within the taxonomy, as evidenced by their lower performance in generating specific occupations compared to major occupations.
@@
9 Acknowledgments
This research / project is supported by the Ministry of Education, Singapore, under its MOE Academic Research Fund Tier 2 programme (Award T2EP20223-0047). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore. The authors would also like to acknowledge support from the Data Science Institute (DSI) at Columbia University and thank the anonymous reviewers for their constructive feedback, which significantly improved the quality of this work.
@@ Appendix A The Lightcast Dataset
Appendix B The Formatting Functions
In our proposed TGRE framework, the formatting function f is a component that converts taxonomic information retrieved for a specific class label y into a grounded rationale r that aligns the LLM reasoning process with the taxonomy. The rationale r serves as an intermediate explanation that clarifies the connection between the input x and the corresponding label y. For example, it explains how a job title-company pair is related to an SOC title and code.
Formally, f takes the retrieved taxonomic information $T(y)$, which consists of information of y, (e.g., occupation description), and transforms it into a rationale $r=f(T(y))$. The rationale is then embedded into in-context examples to guide the reasoning process of LLMs.
In this study, we use a simple implementation of f where the retrieved taxonomic information $T(y)$ is incorporated verbatim in the rationale. This approach minimizes the framework’s complexity while achieving strong empirical results. For example:
Given an input x “records and evidence technician, cobb county police department” and its SOC label y “Police Identification and Records Officers (33-3021.02)” in the validation set, the proposed framework retrieves the matched taxonomic entity $T(y)$ from the O*NET-SOC taxonomy, comprising the followings:
  • Occupation title: Police Identification and Records Officers
  • Occupation code: 33-3021.02
  • Description: Collect evidence at crime scenes, classify and identify fingerprints, and photograph evidence for use in criminal and civil cases.
The verbatim formatting function retrieves the description from $T(y)$ and directly into the grounded rationale r as follows:
  • Grounded rationale: Given the job title ’records and evidence technician’ and company ’cobb county police department,’ the individual is likely expected to collect evidence at crime scenes, classify and identify fingerprints, and photograph evidence for use in criminal and civil cases.
Here, the bolded elements are values derived from the task inputs and taxonomy that replace placeholders within the rationale template.
It is important to note that while we employ a simple verbatim approach in this study, LLMs can be leveraged to make more sophisticated formatting functions. For example, an LLM could rewrite the retrieved description for improved brevity and clarity. Moreover, it could incorporate other taxonomic information beyond descriptions, such as tasks and skills, to generate an enriched summary.
Appendix C Optimization Module in IReRa
The optimization module in IReRa, as depicted in Figure 1 in D’Oosterlinck et al. (2024), employs bootstrapping to iteratively identify the most effective few-shot in-context examples from the validation set. This procedure involves systematically evaluating the performance of various combinations of examples and selecting those that achieve the best results based on a predefined metric. In our experiments, this optimization process identified a specific 1-shot in-context example ($k=1$) for the CoT-based variant in the occupation classification experiment. For the skill classification experiment, the optimization module selected a set of 2-shot in-context examples ($k=2$) for the CoT-based variant. While our TGRE variant also used $k=2$, the examples were randomly chosen, distinguishing them from the optimized examples used by the CoT-based variant.
Appendix D Prompt Templates and In-context Examples
@@ D.1 Knowledge Assessment
D.2 Occupation Classification
This section describes the prompt templates and in-context examples used in the inference and reranking stages of our framework for the occupation classification task. Figures  3 and  4 display TGRE-based and CoT-based prompt templates for occupation classification, respectively. Figures  5 and  6 display k-shot in-context examples for TGRE-based and CoT-based prompt templates for occupation classification, respectively. These in-context examples are integrated into the respective templates by replacing the placeholder \${examples}.
It is important to note that the TGRE prompt templates in Figure 3 do not specify the version of the O*NET taxonomy, unlike the CoT prompt templates in Figure  4. During the prompt engineering, we found that specifying the ONET version did not affect task performance, as LLMs generally lack the precision required to generate SOC titles and codes for specific O*NET-SOC versions. As a result, we opted to omit the version number from TGRE prompts, while keeping them in the CoT prompts as the latter had already been used in prior experiments.
The output format for the CoT-based inference prompt template in Figure 4 is an adaptation from IReRa (D’Oosterlinck et al. 2024), originally designed for multi-label classification. To tailor this format to occupation classification, the output is truncated to retain only the top-1 predicted label.